OpenStudy (anonymous):

Find the product. (m - n + k) (m + 4n - 3k)

OpenStudy (anonymous):

Can someone help out?

OpenStudy (michele_laino):

please you have to apply the distributive poprty of multipication ove addition, namely: \[(m-n+k)(m+4n-3k)=\] \[=m(m+4n-3k)-n(m+4n-3k)+k(m+4n-3k)=...\]

OpenStudy (michele_laino):

please try to complete!

OpenStudy (anonymous):

I did everything in the parentheses but I am having trouble when trying to combine the terms

OpenStudy (michele_laino):

please write all terms you have got

OpenStudy (anonymous):

m^2 +4mn -3km - mn -4n^2 +3kn + km + 4nk - 3k^2

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

now you can add only similar terms, namely: similar terms are terms with identical literal parts. For example: 4mn and -mn are similar terms, because have the same literal part. Now 4mn+(-mn)=3mn do you agree?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok! now another example: -3km and km are similar terms, because, like before, they have the same literal part. So I can add them each other. I will get this: -3km+km=-2km do you agree?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok! now do the same with remaining terms.

OpenStudy (michele_laino):

please

OpenStudy (michele_laino):

of course, write your answer

OpenStudy (anonymous):

m^2 - 4n^2 -3k^2 + 3mn - 2km +7nk

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Cool